Big Python Manual

Release x.y

Your Name Here

April 30, 1999

Organization name, if applicable Street address, if you want to use it

Email: your-email@your.domain

Abstract

Big Python is a special version of Python for users who require larger keys on their keyboards. It accommodates their special needs by ...

CONTENTS

\overline{i} ii

CHAPTER^{ONE} ...^{My chapter.}

\overline{1}

2 APPENDIX

A

...

My appendix. The \appendix markup need not be repeated for additional appendices.

4

MODULE INDEX

6  Module Index

INDEX